OpenStudy (anonymous):

What is the slope of (1,-6) , (9,6)

OpenStudy (anonymous):

have you ever heard of something call the slope formula?

OpenStudy (anonymous):

langrage, be nice

OpenStudy (anonymous):

you would solve this one like the other: 9 - 6 = m(1 - (-6))

OpenStudy (anonymous):

To help you solve in the future, remember that the point slope formula is y - y1 = m ( x - x1)

OpenStudy (anonymous):

or you can use \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{6-(-6)}{9-1}=\frac{12}{8}=\frac{3}{2}\]

OpenStudy (anonymous):

Thankss

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

So the slope is 3/2?

OpenStudy (anonymous):

sure